Sell at Split-Off or Process Further Decision, Alternatives, Relevant Costs

Betram Chemicals Company processes a number of chemical compounds used in producing industrial cleaning products. One compound is decomposed into two chemicals: anderine and dofinol. The cost of processing one batch of compound is $79,000, and the result is 6,000 gallons of anderine and 7,600 gallons of dofinol. Betram Chemicals can sell the anderine at split-off for $11.00 per gallon and the dofinol for $5.95 per gallon. Alternatively, the anderine can be processed further at a cost of $7.80 per gallon (of anderine) into cermine. It takes 4 gallons of anderine for every gallon of cermine. A gallon of cermine sells for $59.

Required:

1. Which alternative is more cost effective and by how much?
NOTE: Do NOT round interim calculations and, if required, round your answer to the nearest dollar.

by $

2. What if the production of anderine into cermine required additional purchasing and quality inspection activity? Every 520 gallons of anderine that undergo further processing require 19 more purchase orders at $12 each and 14 more quality inspection hours at $25 each. Which alternative would be better and by how much?
NOTE: Round interim calculations and your final answer to the nearest cent.
Answer #1

Answer 1

Sale at split-off point is more effective by $ 24,300

Note:

Process further Amount Sale at split-off point Amount
Sales value after processing $      88,500 Sales value $      66,000
Less: Extra Cost $      46,800
Net benefit $      41,700 Net benefit $      66,000

Answer 2

Sale at split-off point is more effective by $ 30,969.23.

Note:

Process further Amount Sale at split-off point Amount
Sales value after processing $ 88,500.00 Sales value $ 66,000.00
Less: Extra Cost $ 46,800.00
Less: Extra Cost of purchasing $   2,630.77
Less: Extra Cost of inspection $   4,038.46
Net benefit $ 35,030.77 Net benefit $ 66,000.00

Calculations:

Process further Amount Sale at split-off point Amount
Sales value after processing =1500*59 Sales value =6000*11
Less: Extra Cost =6000*7.8
Less: Extra Cost of purchasing =6000/520*19*12
Less: Extra Cost of inspection =6000/520*14*25

In case of any doubt, please comment.